# Prompts to Summaries: Zero-Shot Language-Guided Video Summarization

### v1

@@ Abstract
The explosive growth of video data intensified the need for flexible user-controllable summarization tools that can operate without domain-specific training data. Existing methods either rely on datasets, limiting generalization, or cannot incorporate user intent expressed in natural language. We introduce Prompts-to-Summaries; the first zero-shot, text-queryable video-summarizer that converts off-the-shelf video–language models (VidLMs) captions into user-guided skims via large-language-models (LLMs) judging, without the use of training data at all, beating all unsupervised and matching supervised methods. Our pipeline (i) segments raw video footage into coherent scenes, (ii) generates rich scene-level descriptions through a memory-efficient, batch-style VidLM prompting scheme that scales to hours-long videos on a single GPU, (iii) leverages an LLM as a judge to assign scene-level importance scores under a carefully crafted prompt, and finally, (iv) propagates those scores to short segments level via two new metrics—consistency (temporal coherency) and uniqueness (novelty)—yielding fine-grained frame importance. On SumMe and TVSum, our data-free approach surpasses all prior data-hungry unsupervised methods. It also performs competitively on the Query-Focused Video Summarization (QFVS) benchmark, despite using no training data and the competing methods requiring supervised frame-level importance. To spur further research, we release VidSum-Reason (can be found here), a new query-driven dataset featuring long-tailed concepts and multi-step reasoning; our framework attains robust F1 scores and serves as the first challenging baseline. Overall, our results demonstrate that pretrained multimodal models, when orchestrated with principled prompting and score propagation, already provide a powerful foundation for universal, text-queryable video summarization.  We encourage the reader to view our Github page and video.
Figure 1: Prompts to summaries overview. Our zero-shot pipeline turns any video + text query into a tailored highlight reel. We (1) detect scenes, (2) snap boundaries with visual embeddings, (3) caption each scene via a video-language model, and (4) let an LLM rank captions against the query alongside its importance as part of the whole video. Smoothed relevance scores weight every frame, yielding a query-aligned summary—no extra training, ready for personal highlights, generalizes across domains and handles long-tailed, reasoning-heavy queries.
@@ 1 Introduction
The exponential rise in video content, has introduced major challenges in how videos are managed, processed, and consumed [1, 2, 3]. Driven by the ubiquity of smartphones, the explosion of social media, and the expansion of surveillance systems, video are being produced at an unprecedented scale [4], this growth underscores the urgent need for solutions to information overload, escalating storage demands, and the growing difficulty of extracting meaningful insights from raw footage.
1.1 Video summarization
@@
1.2 Limitations of exiting video summarization methods
Despite numerous research efforts [15, 16, 17, 18, 19, 20, 21, 5, 22], current video summarization approaches face two major drawbacks. First, most solutions—whether supervised [18, 19, 20, 21] or unsupervised [15, 16, 17, 5, 22]—depend heavily on the training data, limiting their ability to generalize beyond the distributions they were trained on and thus reducing their real-world applicability. Second, they are constrained to the generic, pre-defined summarization task and cannot accommodate user-specified text queries to guide the summarization process toward specific content or themes. Although there was an effort to generate query-focused video summarization employing a supervised approach [20, 23, 24], the suggested methods rely on dense query-focused user annotations for every frame for training, and they don’t provide support for long tailed queries, or queries requiring reasoning beyond e.g., simple steps object detection .
To be as broadly applicable as possible across diverse video contexts and user needs, summarization algorithms should fulfill two key requirements: (i) they must operate effectively in a zero-shot setting, i.e., without retraining for each new video domain/type, thereby providing a more flexible and generalizable solution; and (ii) they must be text-queryable, giving users the ability to influence the summarization process via natural language prompts—such as “highlight violent scenes”, “remove inappropriate frames,”, and “include expressive emotional moments”.
@@ 1.3 Problem statement
Motivated by the previous discussion, we now define the problem as follows. Given a sequence of video frames
$F=\{f_{t}\}_{t=1}^{T},$
the goal of the traditional video summarization task is to develop an algorithm, denoted as ALG, which receives $F$ as input and produces a video skim composed of selected video segments
$S=\{s_{n}\}_{n=1}^{N},$
ordered chronologically to create a shorter version of the original video. Each segment $s_{n}$ is a contiguous set of consecutive frames, that is,
$s_{n}=\{f_{t_{n}},f_{t_{n}+1},\ldots,f_{t_{n}+\Delta_{n}}\},$
for some starting frame $t_{n}$ and segment length $\Delta_{n}$. This form of summarization aims to maintain temporal continuity, motion, and audio, thereby offering a richer and more natural viewing experience.
@@ 1.4 Main idea
Elsewhere, video language models (VideoLMs) have emerged as a promising tool, enabling language-guided question answering on videos. These models possess some of the desired traits needed to achieve an open-set, text queryable method for video summarization, as they are already trained on a huge amount of internet-scale data. This extensive training allows them to be highly adaptable, making them not confined to a specific data distribution.
Simply put, while there are no VideoLMs specialized for video summarization, VideoLMs can answer textual queries about videos by generating text descriptions (e.g., "What do you see in this scene?"). These text outputs can then be processed by large language models (LLMs), which, using carefully crafted prompts (e.g., "Given this scene description <Here> and the full video summary <Here>, rate the importance of this scene from 1 to 10"), can rank or evaluate the scenes. This flexibility enables this approach to perform summarization tasks across various domains without being limited to one specific application. In addition one can add guidance by text to the prompt, e.g. "Given this scene description <Here>, the full video summary <Here>, and the user query: <Prioritize/highlight/Focus-on scenes with negative emotions>, rate the importance of this scene from 1 to 10".
1.5 Challenges
Notably, its not as easy as described as LLM and VideoLMs where not designed for this purpose. For instance, it is still unclear how to: (i) formulate effective queries for ranking scenes or generating informative text descriptors; (ii) segment videos into semantically meaningful and coherent scenes; (iii) ensure that each scene is both unique and complete; (iv) preserving the temporal context in long videos, especially since video-language models are unable to process the entire video at once due to input length limitations. (v) propagate scene-level importance scores to individual frames—especially since certain frames may better capture the essence of a scene than others; and, most critically, (vi) develop an integrated framework that combines all these stages while leveraging the capabilities of LLMs and VideoLMs in a principled and effective manner. Thus, these challenges and gaps in current knowledge require a careful and critical examination of how to effectively apply VideoLMs and LLMs toward a zero-shot text-queryable video summarization.
@@ 1.6 Our contributions
To this end, we investigate in depth the usage of large-scale pre-trained video-language models with large language models for zero-shot and data free, text-queryable video summarization. Specifically, we explore how VideoLMs and LLMs, can work together as team to semantically align text queries with video content and assess the relevance of individual scenes within the context of the entire video based on textual information, then LLMs can be utilized to assign relevance scores to each video segment based on the textual descriptions generated by the VideoLMs and user-provided queries, as demonstrated in Figure 1. Finally, we delve into methods to propagate scene scoring to frame scores.
With our approach, we provide the following contributions:
  1. (i)
@@
    A novel framework. We provide the first zero-shot text guided video summarization framework that enables generating video summaries conditioned on natural language queries without the need for training data, annotations, or fine-tuning. Our framework can be split into the following stages: (i) Scene detection (ii) Scene description generation (iii) Scene importance within the video, (iv) Frame weight within the scene (v) Frame importance withing the video by aggregating frame weight and scene importance.
  2. (ii)
@@
    State-of-the-art (SOTA) results: Our approach achieves SOTA performance among all unsupervised methods and performs comparably to leading supervised approaches, both of which traditionally require training data. Specifically, our method yields the highest F1-scores on well-established video summarization benchmarks, including the SumMe [25] and TVSum [26] datasets among all unsupervised approaches, and a comparable F1 scores to supervised annotations hungry methods on the query focused video summarization (QFVS) dataset [27]. All of that without relying on any training data.
  3. (iii)
@@
    Experimental study. To determine the optimal design of each component and assess its contribution, we conducted a comprehensive study by systematically omitting and modifying key elements. This analysis quantifies the individual design and impact of each part on the overall performance and reveals insightful findings about the behavior of VideoLMs and LLMs in the context of video summarization.
  4. (iv)
@@
    VidSum-Reason: A new public dataset for text-guided video summarization requiring reasoning. Since the QFVS dataset does not capture long tailed concepts or reasoning based queries, we curated a diverse dataset by selecting videos from the web and annotating each frame according to goal-driven, text-based queries. This allows for evaluating methods beyond standard benchmark datasets. To enable more nuanced analysis, the queries are structured across multiple levels of reasoning, offering a richer testbed for current and future approaches. Indeed, we test our method on this dataset (can be found here), achieving high F1 scores.
@@ 2 Related work
2.1 Standard video summarization
In recent years, a wide range of efforts have focused on generating video skims using deep learning techniques [18, 20]. These approaches can be broadly categorized on the basis of their training paradigms into the following groups.
Supervised methods rely on human-annotated summaries to train models that predict frame- or segment-level importance scores. These approaches often use deep neural networks such as recurrent models or reinforcement learning frameworks to capture temporal dependencies and semantic relevance  [28, 21, 29, 18, 20]. While they achieve strong performance, their reliance on large-scale annotated datasets limits their scalability and generalization across domains. Weakly supervised methods attempt to reduce this dependency by learning from coarser supervision, such as video-level tags, summary-level annotations, or user engagement signals [22, 5]. These methods offer a compromise between label efficiency and learning effectiveness.
@@ 2.1 Standard video summarization
Unsupervised approaches. In contrast, unsupervised approaches aim to generate summaries without any manual annotations. They typically utilize visual features, motion analysis, or statistical cues like diversity and representativeness. Techniques such as clustering, graph modeling, or temporal segmentation (e.g., KTS [30]) are commonly employed to identify informative segments  [17][16][15]. Although unsupervised methods are more scalable, they may struggle to understand and relate temporal relationships between scenes and full videos, high-level semantics, or align with user-specific goals. Notably, they still require training data as they leverage a powerful self-supervised learning method recently adopted for training language models  [17] and incorporating temporal video segmentation with the summarization process  [15].
2.2 Query-focused video summarization
@@
Sharghi et al.[23] introduced the Query-Focused Video Summarization (QFVS) dataset, based on UT Egocentric videos[31], which includes user-defined summaries for a fixed set of concepts. They later proposed a memory network [27] that uses a recurrent attention mechanism to attend to frames and shots based on a user query. However, this approach suffers from limited parallelism and struggles with long-range dependencies—issues our VideoLM-LLM based model addresses. Additionally, their method is restricted to predefined, keyword-based queries from the QFVS dataset. More recently, CLIP-It [20] enabled freeform query input by training on dense video descriptions using language encoders, however it relies on clip, thus limiting its reasoning capabilities to explicitly defined objects rather than long tailed concepts and higher reasoning tasks. Other works [32, 33] also condition summarization on keyword-based queries, however, [32] requires detailed annotations and [32] performance found to be not robust, compared to modern techniques.
@@ 3 Method
We now define our framework, which is designed to effectively capture and represent the essential content of videos without relying on annotated training data. Introduction to the framework. Our proposed pipeline, composed of the following key components, each of which must be carefully evaluated to determine the optimal design (see Figure 1).
@@
    Scene detection: The video is segmented into temporally coherent and semantically meaningful scenes.
@@
    Scene description: For each detected scene, a textual descriptor is generated to summarize its salient content as part of the video in natural language.
@@
    Scene-level importance estimation: Using the generated textual descriptors, an importance score is assigned to each scene, incorporating both semantic relevance and temporal context within the video.
@@
    Frame-level weight estimation within the scene Within each scene, the representativeness of individual frames is evaluated to determine their relative importance (to the scene).
@@
    Fusion of scene and frame importance: For each frame in the full video, the scene-level (scene importance within the video) and frame-level (representativeness within the scene) scores are combined/fused to produce a final frame (importance) score.
Figure 2: Flow diagram for scene boundary detection. Given video frames $\{F_{1},F_{2},\dots,F_{t}\}$ and a threshold $\tau$, scene boundaries are identified based on the mean intensity difference of raw pixels between consecutive frames.
Figure 3: Number of detected scenes across different threshold values for the video Bus in a rock tunnel from the sumMe dataset. At lower thresholds, the number of detected scenes increases, reaching a maximum. As the threshold increases further, fewer scenes are detected due to stricter boundary conditions. The optimal threshold, indicated on the plot, balances over-segmentation and under-segmentation, and is selected based on the point of maximum negative slope.

### v2

# Prompts to Summaries: Zero‑Shot Language‑Guided Video Summarization with Large Language and Video Models

@@ Abstract
The explosive growth of video data intensified the need for flexible user‑controllable summarization tools that operate without training data. Existing methods either rely on domain‑specific datasets, limiting generalization, or cannot incorporate user intent expressed in natural language. We introduce Prompts-to-Summaries: the first zero-shot, text-queryable video-summarizer that converts off-the-shelf video–language models (VidLMs) captions into user-guided skims via large-language-models (LLMs) judging, without the use of training data, beating unsupervised and matching supervised methods. Our pipeline (i) segments video into scenes, (ii) produces scene descriptions with a memory-efficient batch prompting scheme that scales to hours on a single GPU, (iii) scores scene importance with an LLM via tailored prompts, and (iv) propagates scores to frames using new consistency (temporal coherence) and uniqueness (novelty) metrics for fine-grained frame importance. On SumMe and TVSum, our approach surpasses all prior data-hungry unsupervised methods and performs competitively on the Query‑Focused Video Summarization benchmark, where the competing methods require supervised frame-level importance. We release VidSum‑Reason, a query-driven dataset featuring long-tailed concepts and multi-step reasoning, where our framework serves as the first challenging baseline. Overall, we demonstrate that pretrained multi-modal models, when orchestrated with principled prompting and score propagation, provide a powerful foundation for universal, text‑queryable video summarization.
@@
Figure 1: Prompts to summaries overview. Our zero-shot pipeline turns any video + text query into a tailored highlight reel. We (a) detect scenes and snap boundaries with visual embeddings, (b) caption each scene via a video-language model, and (c) let an LLM rank captions against the query alongside its importance as part of the whole video. (d) Smoothed relevance scores weight every frame, yielding a query-aligned summary. No extra training, ready for personal highlights, generalizes across domains and handles long-tailed, reasoning-heavy queries.
@@ 1 Introduction
The exponential rise in video content has introduced major challenges in how videos are managed, processed, and consumed [1, 2, 3]. Driven by the ubiquity of smartphones and the explosion of social media, videos are being produced at an unprecedented scale [4]. This growth underscores the urgent need for solutions to information overload, escalating storage demands, and the increasing difficulty of extracting insights from raw footage.
Video summarization.
@@
Limitations of exiting video summarization methods.
Despite numerous research efforts [15, 16, 17, 18, 19, 20, 21, 5, 22], current video summarization approaches face two major drawbacks. First, most solutions—whether supervised [18, 19, 20, 21] or unsupervised [15, 16, 17, 5, 22]—depend heavily on the training data, limiting their ability to generalize beyond the distributions they were trained on and thus reducing their real-world applicability. Second, they are constrained to the generic, pre-defined summarization task and cannot accommodate user-specified text queries to guide the summarization process toward specific content or themes. Although there was an effort to generate query-focused video summarization employing a supervised approach [20, 23, 24], the suggested methods rely on dense query-focused user annotations for every frame for training, and they don’t provide support for long-tailed queries, or queries requiring reasoning beyond e.g., simple object detection.
To be as broadly applicable as possible across diverse video contexts and user needs, summarization algorithms should fulfill two key requirements: (i) they must operate effectively in a zero-shot setting, i.e., without retraining for each new video domain/type, thereby providing a more flexible and generalizable solution; and (ii) they must be text-queryable, giving users the ability to influence the summarization process via natural language prompts, such as “highlight violent scenes”, “remove inappropriate frames,”, and “include expressive emotional moments”.
@@ 1.1 Problem Statement
Motivated by the previous discussion, we now define the problem as follows. Given a sequence of video frames $F=\{f_{t}\}_{t=1}^{T},$ the goal of the traditional video summarization task is to develop an algorithm, denoted as ALG, which receives $F$ as input and produces a video skim composed of selected video segments $S=\{s_{n}\}_{n=1}^{N},$ ordered chronologically to create a shorter version of the original video. Each segment $s_{n}$ is a contiguous set of consecutive frames, that is, $s_{n}=\{f_{t_{n}},f_{t_{n}+1},\ldots,f_{t_{n}+\Delta_{n}}\},$ for some starting frame $t_{n}$ and segment length $\Delta_{n}$. This form aims to maintain temporal continuity, motion, and audio, thereby offering a richer and more natural viewing experience.
@@ 1.2 Main Idea
Elsewhere, video language models (VideoLMs) have emerged as a promising tool, enabling language-guided question answering on videos. These models possess some of the desired traits needed to achieve an open-set, text-queryable method for video summarization, as they are already trained on a huge amount of internet-scale data. This extensive training allows them to be highly adaptable, making them not confined to a specific data distribution.
Simply put, while there are no VideoLMs specialized for video summarization, VideoLMs can answer textual queries about videos by generating text descriptions (e.g., "What do you see in this scene?"). These text outputs can then be processed by large language models (LLMs), which, using carefully crafted prompts (e.g., "Given this scene description <<Here>> and the full video summary <<Here>>, rate the importance of this scene from 1 to 10"), can rank or evaluate the scenes. This flexibility enables this approach to perform summarization tasks across various domains without being limited to one specific application. In addition one can add guidance by text to the prompt, e.g. "Given this scene description <<Here>>, the full video description <<Here>>, and the user query: <<Prioritize scenes with negative emotions>>, rate the importance of this scene from 1 to 10".
Challenges.
Notably, it’s not as easy as described as LLM and VideoLMs where not designed for this purpose. For instance, it is still unclear how to: (i) formulate effective queries for ranking scenes or generating informative text descriptors; (ii) segment videos into semantically meaningful and coherent scenes; (iii) ensure that each scene is both unique and complete; (iv) preserving the temporal context in long videos, especially since video-language models are unable to process the entire video at once due to input length limitations. (v) propagate scene-level importance scores to individual frames—especially since certain frames may better capture the essence of a scene than others; and, most critically, (vi) develop an integrated framework that combines all these stages while leveraging the capabilities of LLMs and VideoLMs in a principled and effective manner. Thus, these challenges and gaps in current knowledge require a careful and critical examination of how to effectively apply VideoLMs and LLMs toward a zero-shot text-queryable video summarization.
@@ 1.3 Our Contributions
To this end, we investigate in depth the usage of large-scale pre-trained video-language models with large language models for zero-shot and data free, text-queryable video summarization. Specifically, we explore how VideoLMs and LLMs, can work together as team to semantically align text queries with video content and assess the relevance of individual scenes within the context of the entire video based on textual information; LLMs can be utilized to assign relevance scores to each video segment based on the textual descriptions generated by the VideoLMs and user-provided queries, as demonstrated in Figure 1. Finally, we delve into methods to propagate scene scoring to frame scores. With our approach, we provide the following contributions:
  1. 1.
@@
    A novel framework. We provide the first zero-shot text guided video summarization framework that enables generating video summaries conditioned on natural language queries without the need for training data, annotations, or fine-tuning.
  2. 2.
@@
    State-of-the-art (SOTA) results: Our approach achieves SOTA performance among all unsupervised methods and performs comparably to some leading supervised approaches, both of which require domain-related training data. Specifically, our method yields the highest F1-scores on well-established standard video summarization datasets of SumMe and TVSum, among all unsupervised approaches, and comparable F1 scores to supervised (per-frame) annotations hungry methods on the query-focused video summarization (QFVS) dataset. All of that without relying on any training data.
  3. 3.
@@
    Experimental study. To determine the optimal design of each component and assess its contribution to our pipeline, we conducted a comprehensive study by systematically omitting and modifying key elements. This analysis quantifies the individual design and impact of each part on the overall performance and reveals insightful findings about the behavior of VideoLMs and LLMs in the context of video summarization.
  4. 4.
@@
    VidSum‑Reason: A new public dataset for text-guided video summarization requiring reasoning. Since the QFVS dataset does not capture long-tailed concepts or reasoning-based queries, we curated a diverse dataset by selecting videos from the web and annotating each frame according to goal-driven, text-based queries. This allows for evaluating methods beyond standard benchmark datasets. To enable more nuanced analysis, the queries are structured across multiple levels of reasoning, offering a richer testbed for current and future approaches. We test our method on this dataset, achieving high F1 scores.
Novelty of our work. To enable our training-free, text-guided video summarization framework, we introduce four main novelties: (i) formulating video summarization as LLM-based importance scoring over structured caption units, (ii) scaling video description generation of VideoLM to long videos via memory-aware batched description generation, (iii) designing a reproducible, rubric-driven prompting protocol for stable importance judgments, and (iv) proposing lightweight score propagation to derive frame-level saliency—all in a fully zero-shot, domain-agnostic setting (see Appendix Section F for more details).
@@ 2 Related work
Standard video summarization.
@@
In recent years, a wide range of efforts have focused on generating video skims using deep learning techniques [18, 20]. These approaches can be broadly categorized on the basis of their training paradigms into the following groups. Supervised methods rely on human-annotated summaries to train models that predict frame- or segment-level importance scores. These approaches often use deep neural networks such as recurrent models or reinforcement learning frameworks to capture temporal dependencies and semantic relevance  [25, 21, 26, 18, 20]. While they achieve strong performance, their reliance on large-scale annotated datasets limits their scalability and generalization across domains. Weakly supervised methods attempt to reduce this dependency by learning from coarser supervision, such as video-level tags, summary-level annotations, or user engagement signals [22, 5]. These methods offer a compromise between label efficiency and learning effectiveness. Unsupervised approaches. In contrast, unsupervised approaches aim to generate summaries without any manual annotations. They typically utilize visual features, motion analysis, or statistical cues like diversity and representativeness. Techniques such as clustering, graph modeling, or temporal segmentation (e.g., KTS [27]) are commonly employed to identify informative segments  [17][16][15]. Although unsupervised methods are more scalable, they may struggle to understand and relate temporal relationships between scenes and full videos, high-level semantics, or align with user-specific goals. Notably, they still require training data as they leverage a powerful self-supervised learning method adopted for training language models  [17] and incorporating temporal video segmentation with the summarization process  [15].
Recent zero- and few-shot approaches have explored video summarization with minimal supervision. Contrastive Zero-Shot [28] ranks frames or segments based on pre-trained video representations learned through a contrastive objective, without relying on labeled summaries. Few-shot methods, such as STeMI [29], rely on a limited set of annotated examples to learn segment importance. Both approaches, however, do not support text-queryable summarization and cannot directly generate summaries conditioned on arbitrary user queries.
Query-focused video summarization.
@@
[23] introduced the Query-Focused Video Summarization (QFVS) dataset, based on UT Egocentric videos [30], which includes user-defined summaries for a fixed set of concepts. They later proposed a memory network [31] that uses a recurrent attention mechanism to attend to frames and shots based on a user query. However, this approach suffers from limited parallelism and struggles with long-range dependencies—issues our VideoLM-LLM based model addresses. Additionally, their method is restricted to predefined, keyword-based queries from the QFVS dataset. More recently, CLIP-It [20] enabled freeform query input by training on dense video descriptions using language encoders, however it relies on clip, thus limiting its reasoning capabilities to explicitly defined objects rather than long tailed concepts and higher reasoning tasks. Other works [32, 33] also condition summarization on keyword-based queries, however, [32] requires detailed annotations and [32] performance found to be not robust, compared to modern techniques.
Limited Supervision and Temporal Reasoning.
Several recent methods tackle video–text alignment under limited supervision.  [34] leverages memory modules to improve semantic generalization with scarce annotations, illustrating one approach to limited supervision in multi‑modal settings.  [35] constructs a hierarchical segment tree from a single labeled frame and introduces self‑supervised training strategies to score candidate segments under extreme label scarcity. While both methods focus on localizing query‑relevant temporal segments, our approach differs by leveraging pretrained  VideoLM and LLM prompting to score semantic importance without any training data.  [36] highlights another recent direction in video–text reasoning that could potentially be explored for future fine‑grained scoring in summarization. In contrast, our framework emphasizes zero‑shot, text‑guided summarization by combining scene descriptions and LLM‑driven scoring to produce semantically relevant skims across diverse domains.
@@ 3 Method
We now define our framework, which is designed to effectively capture and represent the essential content of videos without relying on annotated training data.
Introduction to the framework.
Our proposed pipeline (see Figure 1), composed of the following key components, each of which must be carefully evaluated to determine the optimal design.
@@
    Scene detection: The video is segmented into temporally coherent and semantically meaningful scenes (Figure 1a).
@@
    Scene description: For each detected scene, a textual descriptor is generated to summarize its salient content as part of the video in natural language (Figure 1b).
@@
    Scene-level importance estimation: Using the generated textual descriptors, an importance score is assigned to each scene, incorporating both semantic relevance and temporal context within the video (Figure 1c).
@@
    Frame-level weight estimation within the scene Within each scene, the representativeness of individual frames is evaluated to determine their relative importance to the scene (Figure 1d).
@@
    Fusion of scene and frame importance: For each frame in the full video, the scene-level (Step 3: scene importance within the video) and frame-level (Step 4: representativeness within the scene) scores are combined/fused to produce a final frame (importance) score (Figure 1d).